find the slope-intercept equation for a line when given m=2/9 and passes through (3,7)
Answer by mananth(16946)

the general equation in point slope form is
y-y1=m(x-x1)
m= slope, (x1,y1)=>point
y-7=2/9 (x-3)
y= (2/9)x-(2/9)(-3)+7
y=(2/9)x -2/3+7
y=(2/9)x+19/3